A body travel 200cm in the first two seconds and 220cm in the next 4 seconds with deceleration. Find the velocity at the end of 7th second 

Grade:

2 Answers

shubham sharma
10 years ago

dear , just make two equa.

using S=ut plus ½at''2

1st , S=2m, t=2s and with -a

For nxt S=4.4m , t=6s solve these and just use 1st law of motion to find vel. After 7th sec.

Cherry
6 years ago
 
 
S=ut + 1/2 at2  
200=2u +2a
420=6u +18a
by solving these two , a = -15, by substituting this any one of the equations, u = 115
for 7 seconds V= U + at
                       V = 115 – 15×7
                        V = 115-105 =10